OpenStudy (anonymous):

What will be the angle between two plane mirrors if 8 images of an object placed symmetrically are to be formed ?

OpenStudy (anonymous):

Hi s.dhanush Let's Imagine this. If we put one miror we have two images right? the real object and one image in the mirror.  If we have two plane mirros at 90 deg then we have 4 images. Why is that? the second mirror has an image of itself in the first one and vise versa  so we have 4 images, one real and 3 Imaginary Image. and u can see that our wold is divided to four "copies" of itself. thats because 360/90 =4 so now if u want 8 images u just have to make 7 "copies" thats 360/8 = 45 or if u want 8 Imaginary Images u need 9 "worlds" thats 360/9=40 good luck! :) Or

OpenStudy (anonymous):

360/angle -1 =images= a=40